Rewrite and solve the word problems as equations.
The length of a rectangle is 1m longer than the breadth.
If the perimeter is 22m, what is the length.

Answer by JulietG(1812)

Perimeter = 2 length + 2 width
22 = 2L + 2W
L = W + 1
22 = 2(W+1) + 2W
22 = 2W + 2 + 2W
22 = 4W + 2
20 = 4W
5 = W
If the width is 5, then the length is 6.
5 + 5 + 6 + 6 = 22